a mother is 24 years older than her daughter if the daughter's age is x year:
(a) Express the mothers age in terms of x?
(b) Find x when the daughter age is one-third of her mother's age

Answer by t0hierry(194)

mother age = m = daughter age + 24 = x + 24
m = x + 24
m = 3x
Now you have to solve
m = x + 24
hence
3x = x + 24
2x = 24
x = 12
check
m = x + 24 = 36 = 3*12.